the area of a circle inscribed in a right triangle with sides of length 8 metre, 15 metre and 17 metre is
Answer by richard1234(7193)

Area = (1/2)(8 m)(15 m) = 60 m^2

Also, Area = (inradius)*(semi-perimeter), so

60 m^2 = (inradius)*(20 m) --> inradius = 3 m. The area of the inscribed circle is 9pi m^2.